40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16860     Entry time: Tue May 17 18:43:38 2022     In reply to: 16859     Reply to this: 16861   16862   16863
Author: Anchal 
Type: Update 
Category: BHD 
Subject: Placed SRM in ITMY Chamber 

[Anchal, Paco, Yuta]

SRM Placement

  • SRM was moved from its parked location to the nominal position in the ITMY chamber.
  • This imbalanced the table a lot as all SOS towers ended up on the south side of the table.
  • I needed additionally three SOS tower side walls to recover the balance of the table.
    • I initially tried to use a level meter on my phone which claimed to have 0.1 degrees of accuracy. But it turned out to be a bad idea.
    • Eventually, I used the spirit bubble level meter we have, along with the OSEM values of ITMY, AS1, and AS4.
  • At the end, the table is balanced as it was before, all SOS are damping usually.

SRM Sat Amp Box setup

  • SRM Gold Box Sat Amp was found near the BS chamber.
  • This box was moved to the ITMY chamber.
  • The new flange on the East end was marked earlier for SRM. This flange on the vacuum side was connected with new in-vacuum blue ribbon cables.
  • We had previously moved the cable post for SRM (40m/16849) behind AS2. This cable post is connected to the old in-vacuum cable.
  • It would have changed the table balance to remove this cable post and connect new in-vacuum cables to it, so we decided to do this in the next vent when we put the BHD board on the table.
  • For now, we connected the old in-vacuum cable to the new in-vacuum blue ribbon cables inside.
    • Note, that the old in-vacuum cable has a gender flipping section which also mirrors the pin layout.
    • We installed pin mirroring cables on the outside between the Sat Amp Box and the vacuum flange to revert back the additional mirroring.
    • However it happened, now the Sat Amp Box is working, with all OSEMs and coils alive.
  • One peculiarity we found was that the SRM face OSEMs have only about 250-300 um of range, which is roughly 3 times less than the other OSEMs in other SOSs.
  • SRM side OSEM however behaves normally.
  • After installment, at the free-hanging state, SRM LL OSEM is saturated (too bright) and other face OSEMs are close to total brightness state.
  • We'll first put the alignment offsets to get the SRM perpendicular to the beam coming from SR2 and then center the OSEMs in this tiny range.
  • The low OSEM range could be due to improper biasing from the Sat Amp Box. Hopefully, with new electronics, this issue would go away in future.
ELOG V3.1.3-